The base of the parallelogram is 154 cm and its diagonal make an angle 28° and 43° with the base. Find the length of the longer diagonal. 


1
Expert's answer
2021-10-19T14:05:32-0400

we have triangle ABC,

where AB = 154 cm is base,

BC, AC are halves of diagonals,

"\angle BAC=\alpha=28\degree,\angle ABC=\beta=43\degree"

"\angle ACB=\gamma=180\degree-28\degree-43\degree=109\degree"


Then:

"\frac{sin\alpha}{BC}=\frac{sin\beta}{AC}=\frac{sin\gamma}{AB}"


"AC=ABsin\beta/sin\gamma=154sin43\degree/sin109\degree=111.08" cm

"BC=ABsin\alpha/sin\gamma=154sin28\degree/sin109\degree=76.46" cm


So, the length of the longer diagonal:

"2AC=2\cdot111.08=222.16" cm